a piece of wood is floating at the surface of some water as illustrated. the wood has a circular cross section and a height h=3.0cm. the density of the wood is 0.41g/cm^3. what is the distance y from the surface of the water to the bottom of the wood?

the water to the battom of the wood? h Wood
Answer #1

height h=3.0cm.

the density of the wood= 0.41g/cm3

We need to find the distance y from the surface of the water to the bottom of the wood

We know that density=mass/volume

0.41 g/cm3*h*area=distance*area

Distance=0.41*3=1.2 cm